# STAR-GO: Improving Protein Function Prediction by Learning to Hierarchically Integrate Ontology-Informed Semantic Embeddings

### v1

@@ 3.3 Hierarchical ordering and ontology-informed embeddings improve performance
Removing either the textual or structural component from the GO embeddings results in performance degradation, most notably for BP, where Macro AUPR drops from $0.288$ to $0.267$ (structural-only) or $0.195$ (textual-only). The textual-only variant performed substantially worse here than in the zero-shot setting, where it achieved the highest AUCs. This discrepancy suggests that textual embeddings (SBERT-BioBERT) are effective for transferring to unseen terms but less so when trained end-to-end, whereas structural embeddings (anc2vec) provide a stronger inductive bias and stability when full supervision is available. The STAR-GO performed the best across ontologies leveraging textual and structural information.
We further compared different strategies for integrating GO embeddings with protein representations. In the decoder-based configuration, GO embeddings attend to each other with a causal mask such that GO terms are processed in hierarchical or GO ID-based order. The encoder-based configuration, which attends GO term sequence embeddings bidirectionally without ordered decoding, achieved competitive AUCs in CC and MF (0.990 and 0.995) and the highest CC $F_{\text{max}}$ ($0.667$), but could not be evaluated on BP due to memory constraints. The MLP integration, which pairs mean-pooled protein embeddings with individual GO term embeddings, produced similarly high AUCs (up to 0.995) but underperformed in $F_{\text{max}}$, Macro AUPR, and Micro AUPR compared to the decoder variant with GO hierarchical ordering.
@@ 5 Competing interests
6 Author contributions statement
@@ Appendix A Supplementary Materials
Table 4 summarizes the number of protein sequences and annotated GO terms in the training, validation, and test splits used for model development. Table 5 presents the corresponding statistics for the dataset employed in zero-shot evaluation, including per-subontology protein and term counts.
Table 4: Dataset statistics for training, validation and test splits. The number of protein sequences in each split and the number of GO terms with annotations per ontology.
@@
Table 5: Statistics for UniProtKB–SwissProt 2021_4 dataset used in zero-shot evaluation. Per-subontology protein counts in each split and term count is shown.
@@
Zero-shot capable methods. DeepGOZero [12] pioneered zero-shot protein function prediction by representing GO terms as geometric n-balls that satisfy formal Description Logic axioms, operating on InterPro domain features rather than sequence embeddings. DeepGO-SE [14] extends this geometric approach through approximate semantic entailment across multiple models while incorporating ESM2 protein language models. Both represent the current state-of-the-art in zero-shot prediction.
@@
We compared our GO embedding module against several baseline GO embedding techniques on the subsumption prediction task (Table 6), including OWL2Vec* [4], anc2Vec [5], and GT2Vec [27]. To assess the contribution of semantic information, we also evaluated pretrained language models that encode GO term definitions without structural supervision. Specifically, we included two task-specific BERT models: BioBERT [15], pretrained on large-scale biomedical text, and SBERT-BioBERT [19], which fine-tunes BioBERT on sentence similarity tasks using the SentenceTransformers framework.
@@
As shown in Table 6, anc2vec achieved the highest accuracy in recovering subsumption relations, consistent with its explicit optimization for structural hierarchy. Sentence-based embeddings performed worse overall, though fine-tuned semantic embeddings (SBERT-BioBERT) provided a substantial improvement over non-fine-tuned BioBERT, demonstrating the effectiveness of fine-tuning the SBERT-BioBERT embeddings. Our GO embedding module, which integrates semantic embeddings with structural reconstruction objectives, performed comparably to OWL2Vec* across metrics and surpassed purely semantic models. Despite anc2vec’s stronger subsumption scores, our ablation study (main text, Table 3) showed that these embeddings generalized less effectively to downstream protein function prediction. In contrast, our module’s semantic–structural integration improved transfer to function prediction, confirming the value of including semantic information alongside hierarchical structure.
Table 6: Performance of GO embedding techniques. anc2vec and OWL2Vec* are the same as in their respective papers. BioBERT uses mean-pooled sentence embeddings of term descriptions. SBERT-BioBERT is a BioBERT model fine-tuned for sentence similarity tasks. Our GO embedding method fine-tunes SBERT-BioBERT further with structure recovery objectives. We evaluated on two editions of GO: full and basic.
Hits@1 Hits@10 Hits@5 MRR
GO Edition Method

### v2

@@ 3.3 Hierarchical ordering and ontology-informed embeddings improve performance
Removing either the textual or structural component from the GO embeddings results in performance degradation, most notably for BP, where Macro AUPR drops from $0.288$ to $0.267$ (structural-only) or $0.195$ (textual-only). The textual-only variant performed substantially worse here than in the zero-shot setting, where it achieved the highest AUCs. This discrepancy suggests that textual embeddings (SBERT-BioBERT) are effective for transferring to unseen terms but less so when trained end-to-end, whereas structural embeddings (anc2vec) provide a stronger inductive bias and stability when full supervision is available. STAR-GO performed the best across ontologies leveraging textual and structural information.
We further compared different strategies for integrating GO embeddings with protein representations. In the decoder-based configuration, GO embeddings attend to each other with a causal mask such that GO terms are processed in hierarchical or GO ID-based order. The encoder-based configuration, which attends GO term sequence embeddings bidirectionally without ordered decoding, achieved competitive AUCs in CC and MF (0.990 and 0.995) and the highest CC $F_{\text{max}}$ ($0.667$), but could not be evaluated on BP due to memory constraints. The MLP integration, which pairs mean-pooled protein embeddings with individual GO term embeddings, produced similarly high AUCs (up to 0.995) but underperformed in $F_{\text{max}}$, Macro AUPR, and Micro AUPR compared to the decoder variant with GO hierarchical ordering. Additionally, we evaluated ESM-1 and ESM-2 protein embeddings and found that ProtT5 performed better for our model (see Supplementary Table S4).
3.4 STAR-GO discovers key residues in the zero-shot setting
The decoder cross-attention weights indicate protein residues the model considers relevant to each GO term. To aggregate the signals across heads and layers, we applied recursive averaging to cross-attention matrices, inspired by attention rollout [1]. Let $\bar{\mathbf{A}}^{(l)}=\frac{1}{H}\sum_{h=1}^{H}\mathbf{A}^{(l)}_{h}\in\mathbb{R}^{T\times L}$ denote the head-averaged cross-attention at decoder layer $l$. The rollout is computed recursively as:
$\mathbf{R}^{(1)}$ $=\bar{\mathbf{A}}^{(1)}$
$\mathbf{R}^{(l)}$ $=\tfrac{1}{2}\bigl(\bar{\mathbf{A}}^{(l)}+\mathbf{R}^{(l-1)}\bigr),\quad l=2,\ldots,N_{\text{dec}},$ (4)
equally weighting the current layer’s attention and the accumulated score. Row $t$ of the final rollout $\mathbf{R}=\mathbf{R}^{(N_{\text{dec}})}\in\mathbb{R}^{T\times L}$ yields the per-residue attention profile $\mathbf{r}_{t}\in\mathbb{R}^{L}$ for GO term $t$. We applied this procedure to STAR-GO in the zero-shot setting to determine whether cross-attention captures biologically meaningful signals for a function the model has never seen. We selected two held-out Molecular Function terms related to DNA-binding transcription factor activity: GO:0001228 (transcription activator, zero-shot AUC = 0.938) and GO:0001227 (transcription repressor, zero-shot AUC = 0.891). We computed attention rollout scores for three PDB chains: (GO:0001228) 4AWL chain B (NF-YB), 2F8X chain C (RBPJ/CSL), and (GO:0001227) 2HDC chain A (FOXD3). As ground truth for DNA-binding residues, we obtained experimentally determined DNA-contact sites from the BioLiP database [29].
4AWL-B
Figure 2: Interpretability analysis of STAR-GO’s zero-shot cross-attention for two held-out DNA-binding transcription factor terms: GO:0001228 (activator) and GO:0001227 (repressor). (a) Attention rollout and binding-site residues for 4AWL-B. Above: per-residue attention rollout scores with DNA-contact sites (red). Below: heatmap of per-residue attention averaged over heads for each decoder layer. (b) 3-D structure of 4AWL-B coloured by attention rollout score (blue $\to$ red indicates increasing attention), with DNA shown in green. (c) ROC curves evaluating agreement between attention-derived residue scores and BioLiP binding-site annotations.
Figure 2a shows the per-residue attention rollout profile for 4AWL-B, with head-averaged attention at each decoder layer as a heatmap and the aggregated rollout score plotted above. The BioLiP-annotated DNA-contact residues (red dots) coincide with regions of elevated attention, particularly around residues 25-30, corresponding to the NF-YB DNA-contact region, with a strong signal in the early decoder layers. The 3D structure (Figure 2b) confirms that residues receiving high rollout scores (red) are spatially proximal to the bound DNA (green). Figure 2c reports ROC curves for all three chains: 4AWL-B achieves an AUROC of 0.841, 2HDC-A 0.789, and 2F8X-C 0.727, all substantially above the random baseline.
Critically, both GO:0001228 and GO:0001227 were excluded from all protein-term associations during training; STAR-GO has never observed any protein annotated with either term. STAR-GO’s elevated attention to experimentally validated DNA-contact residues arises solely from the information in the term’s embedding and its alignment with sequence features learned through supervision on other GO terms. This demonstrates that the decoder learns a generalizable mapping between GO term semantics and residue-level features that extends beyond the training vocabulary.
@@ 5 Conclusion
Author contributions statement
@@ Appendix A Supplementary Materials
Table S4 summarizes the number of protein sequences and annotated GO terms in the training, validation, and test splits used for model development. Table S5 presents the corresponding statistics for the dataset employed in zero-shot evaluation, including per-subontology protein and term counts.
Table S4: Dataset statistics for training, validation and test splits. The number of protein sequences in each split and the number of GO terms with annotations per ontology.
@@
Table S5: Statistics for UniProtKB–SwissProt 2021_4 dataset used in zero-shot evaluation. Per-subontology protein counts in each split and term count is shown.
@@
Transformer Module
The encoder and decoder of STAR-GO follow the standard pre-norm Transformer architecture [26] with multi-head attention (MHA), position-wise feed-forward networks (FFN), and layer normalization (LN). Below we provide the full equations for completeness.
Encoder Self-Attention
Let $\mathbf{H}^{(l)}_{\text{enc}}\in\mathbb{R}^{L\times d}$ denote the hidden representation of the protein sequence at the $l$-th encoder layer, where $\mathbf{H}^{(0)}_{\text{enc}}$ is the projected input (see main text). Each layer updates these representations as follows:
where $\mathrm{MHA}$ denotes multi-head attention, $\mathrm{FFN}$ is a position-wise feed-forward network, $\mathrm{LN}$ denotes layer normalization, and $\tilde{\mathbf{M}}_{\text{seq}}$ is an additive mask for sequence padding. After $N_{\text{enc}}$ layers, the encoder outputs refined residue embeddings $\mathbf{H}^{(N_{\text{enc}})}_{\text{enc}}$.
Decoder Self-Attention and Cross-Attention
Let $\mathbf{H}^{(l)}_{\text{dec}}\in\mathbb{R}^{T\times d}$ denote the hidden representation of the GO terms at decoder layer $l$, where $\mathbf{H}^{(0)}_{\text{dec}}$ is the projected GO embedding input (see main text). Each decoder layer is computed as:
$\textbf{Self-attention:}\;\hat{\mathbf{H}}^{(l)}_{\text{dec}}$ $=\mathrm{MHA}\!\left(\mathbf{H}^{(l-1)}_{\text{dec}},\mathbf{H}^{(l-1)}_{\text{dec}};\tilde{\mathbf{M}}_{\text{dec}}\right),$ (7)
$\textbf{Cross-attention:}\;\tilde{\mathbf{H}}^{(l)}_{\text{dec}}$ $=\mathrm{MHA}\!\left(\mathbf{S}^{(l)}_{\text{dec}},\mathbf{H}^{(N_{\text{enc}})}_{\text{enc}};\mathbf{0}\right),$ (9)
$\textbf{Feed-forward:}\;\mathbf{H}^{(l)}_{\text{dec}}$ $=\mathrm{LN}\!\left(\mathbf{C}^{(l)}_{\text{dec}}+\mathrm{FFN}\!\left(\mathbf{C}^{(l)}_{\text{dec}}\right)\right)$. (11)
The decoder’s self-attention mask $\tilde{\mathbf{M}}_{\text{dec}}$ is a lower-triangular matrix that respects the topological ordering of GO terms, ensuring that each term attends only to its ancestors and preceding terms in the hierarchy. Cross-attention uses no mask ($\mathbf{0}$), allowing each GO term to attend to all encoded protein residues.
Prediction Head and Training Objective
For each GO term $t$, the final decoder output is mapped through a two-layer feed-forward projection:
where $\hat{y}_{t}$ denotes the predicted probability that the protein is associated with GO term $t$, and $\sigma(\cdot)$ is the sigmoid activation function.
Given binary ground-truth $\mathbf{y}\in\{0,1\}^{T}$, where $y_{t}=1$ if the protein is annotated with GO term $t$ and 0 otherwise, the model is trained using binary cross-entropy:
@@
Zero-shot capable methods. DeepGOZero [14] pioneered zero-shot protein function prediction by representing GO terms as geometric n-balls that satisfy formal Description Logic axioms, operating on InterPro domain features rather than sequence embeddings. DeepGO-SE [16] extends this geometric approach through approximate semantic entailment across multiple models while incorporating ESM2 protein language models. More recently, TransFew [4] follows a method similar to STAR-GO by combining BioBert embeddings of GO definitions with GO hierarchy through a GCN, and cross-attends them with refined ESM embeddings.
@@
We compared our GO embedding module against several baseline GO embedding techniques on the subsumption prediction task (Table S6), including OWL2Vec* [6], anc2Vec [7], and GT2Vec [31]. To assess the contribution of semantic information, we also evaluated pretrained language models that encode GO term definitions without structural supervision. Specifically, we included two task-specific BERT models: BioBERT [17], pretrained on large-scale biomedical text, and SBERT-BioBERT [22], which fine-tunes BioBERT on sentence similarity tasks using the SentenceTransformers framework.
@@
As shown in Table S6, anc2vec achieved the highest accuracy in recovering subsumption relations, consistent with its explicit optimization for structural hierarchy. Sentence-based embeddings performed worse overall, though fine-tuned semantic embeddings (SBERT-BioBERT) provided a substantial improvement over non-fine-tuned BioBERT, demonstrating the effectiveness of fine-tuning the SBERT-BioBERT embeddings. Our GO embedding module, which integrates semantic embeddings with structural reconstruction objectives, performed comparably to OWL2Vec* across metrics and surpassed purely semantic models. Despite anc2vec’s stronger subsumption scores, our ablation study (main text, Table 3) showed that these embeddings generalized less effectively to downstream protein function prediction. In contrast, our module’s semantic–structural integration improved transfer to function prediction, confirming the value of including semantic information alongside hierarchical structure.
Table S6: Performance of GO embedding techniques. anc2vec and OWL2Vec* are the same as in their respective papers. BioBERT uses mean-pooled sentence embeddings of term descriptions. SBERT-BioBERT is a BioBERT model fine-tuned for sentence similarity tasks. Our GO embedding method fine-tunes SBERT-BioBERT further with structure recovery objectives. We evaluated on two editions of GO: full and basic.
GO Edition Method Hits@1 Hits@10 Hits@5 MRR
@@
Impact of residue embeddings across different protein language models
To assess the impact of the protein language model on downstream function prediction, we compared three residue-level embedding methods within STAR-GO’s architecture: ESM-1b [23], ESM-2 [19], and ProtT5 [8]. All other model components, including the GO embedding module, decoder architecture, and training procedure, were held constant; only the frozen residue embeddings provided to the encoder were varied. Table S7 reports results under our standard supervised evaluation protocol across all three ontologies.
ProtT5 embeddings consistently outperform both ESM variants across most metrics. The improvement is most pronounced for Molecular Function, where ProtT5 achieves an $F_{\text{max}}$ of $0.719$ compared to $0.675$ for both ESM-1b and ESM-2, and a Macro AUPR of $0.620$ versus 0.5740.5740.5760.576. Similar gains are observed for Biological Process and Cellular Component in Macro AUPR and Micro AUPR. AUC scores are comparable across all three embeddings, indicating that the discriminative advantage of ProtT5 is most evident in precision-recall–based metrics where ranking quality among positive terms matters most. These results motivated our adoption of ProtT5 as the default protein encoder in STAR-GO.
Table S7: Comparison of protein language model embeddings for protein function prediction. We evaluate three residue-level embedding methods: ESM1b, ESM2, and ProtT5. Results are reported across Biological Process (BP), Cellular Component ($C$), and Molecular Function (MF) ontologies. ProtT5 embeddings (used in STAR-GO) consistently outperform ESM variants across most metrics, with particularly notable gains in Macro AUPR and Fmax.
Embeddings Macro AUPR Micro AUPR AUC Fmax
Residue Embeddings BP $C$ MF BP $C$ MF BP $C$ MF BP $C$ MF
ESM1b 0.253 0.329 0.574 0.321 0.408 0.631 0.980 0.988 0.995 0.539 0.639 0.675
ESM2 0.268 0.349 0.576 0.321 0.406 0.628 0.975 0.985 0.994 0.544 0.640 0.675
ProtT5 0.288 0.379 0.620 0.351 0.455 0.675 0.989 0.988 0.995 0.548 0.659 0.719
Zero-shot experiment sensitivity analysis
To assess the stability of STAR-GO’s zero-shot predictions with respect to random initialization, we repeated the zero-shot evaluation across five independent training seeds, keeping all hyperparameters fixed. Tables S8 and S9 report per-term AUC scores for each seed along with the mean and standard deviation.
Across all 16 held-out GO terms, the model exhibits low variance in most terms, with most standard deviations below $0.03$. Cellular Component terms show particularly stable predictions, exemplified by GO:0005762 (mean AUC $0.9961\pm 0.0016$). Overall, the consistently low standard deviations confirm that the zero-shot generalization is robust and not an artifact of a particular random seed.
Table S8: Zero-shot protein function prediction AUC scores (Biological Process) across 5 seeds.
Biological Process
Seed GO:0000381 GO:0032729 GO:0032755 GO:0032760 GO:0046330 GO:0051897 GO:0120162
1 0.9735 0.8936 0.9042 0.9064 0.9685 0.9073 0.7754
2 0.9831 0.9201 0.9152 0.9630 0.9006 0.9034 0.8103
3 0.9709 0.8891 0.9030 0.9254 0.9609 0.9347 0.8290
4 0.9715 0.9058 0.9070 0.8935 0.9281 0.9223 0.8051
5 0.9849 0.9106 0.9162 0.9132 0.9556 0.9341 0.8115
Mean $\pm$ Std $0.9768\pm 0.0067$ $038\pm 0.0126 0.909$$091\pm 0.0062$ $0.9203\pm 0.0265$ $0.9427\pm 0.0281$ $0.9203\pm 0.0146$ $0.8063\pm 0.0195$
Table S9: Zero-shot protein function prediction AUC scores (Cellular Component & Molecular Function) across 5 seeds.
Cellular Component Molecular Function
Seed GO:0005762 GO:0022625 GO:0042788 GO:1904813 GO:0001227 GO:0001228 GO:0003735 GO:0004867 GO:0005096
1 0.9941 0.9672 0.8809 0.7675 0.8956 0.9271 0.7650 0.6246 0.9306
2 0.9948 0.9216 0.9056 0.7231 0.9270 0.9476 0.9219 0.8552 0.9367
3 0.9973 0.9746 0.9625 0.7587 0.9476 0.9526 0.6933 0.9030 0.9243
4 0.9975 0.9622 0.9254 0.6598 0.9387 0.9413 0.8459 0.7896 0.9405
5 0.9969 0.8599 0.9041 0.7939 0.9253 0.9442 0.9375 0.7253 0.9184
Mean $\pm$ Std $0.9961\pm 0.0016$ $0.9371\pm 0.0478$ $0.9157\pm 0.0305$ $0.7406\pm 0.0518$ $0.9269\pm 0.0197$ $0.9426\pm 0.0096$ $0.8327\pm 0.1039$ $0.7796\pm 0.1096$ $0.9301\pm 0.0090$